In the ground state of the hydrogen atom the electron describe a circular orbit of radius 5.3×10^-11m at a speed 2.2×10^6ms^-1(a) what is the average electric current resulting from the electron movement (b) calculate the value of B established by the electron at the nuclues


1
Expert's answer
2021-06-30T10:18:54-0400

a)

"T=\frac{2\pi r}{v},"

"I=\frac eT=\frac{ev}{2\pi r}=1.06~mA,"

b)

"\frac{mv^2}r=evB,"

"B=\frac{mv}{er}=236~kT."